# How Good is Your Wikipedia?

### v1

@@ Abstract
Wikipedia’s perceived high quality and broad language coverage have established it as a fundamental resource in multilingual NLP. In the context of low-resource languages, however, these quality assumptions are increasingly being scrutinised. This paper critically examines the data quality of Wikipedia in a non-English setting by subjecting it to various quality filtering techniques, revealing widespread issues such as a high percentage of one-line articles and duplicate articles. We evaluate the downstream impact of quality filtering on Wikipedia and find that data quality pruning is an effective means for resource-efficient training without hurting performance, especially for low-resource languages. Moreover, we advocate for a shift in perspective from seeking a general definition of data quality towards a more language- and task-specific one. Ultimately, we aim for this study to serve as a guide to using Wikipedia for pretraining in a multilingual setting. We also introduce a package that implements our data cleaning, tokenization, and model training pipeline. [^1] https://github.com/akulmizev/texieve
@@ 1 Introduction
Wikipedia is an invaluable resource in the field of NLP. Much of its usefulness stems from its open-source nature, through which a community of authors, editors, and administrators can collaborate on the documentation of an ever-changing world. Such a framework ensures that information contained therein is not only current, but also well-written, accurate, and relevant to any given subject. For these reasons, Wikipedia has been employed for a myriad of NLP applications, ranging from language model training and benchmarking (Merity et al., 2016), question-answering (Hewlett et al., 2016; Rajpurkar et al., 2016; Joshi et al., 2017) and knowledge-base creation (Vrandečić and Krötzsch, 2014; Lehmann et al., 2024), among many others. Wikipedia is a staple resource in multilingual NLP, where it is commonly employed for pretraining multilingual language models (LMs) (Devlin et al., 2019; Conneau and Lample, 2019) and creating benchmarks for downstream tasks, such as named-entity recognition Rahimi et al. (2019) and machine translation Schwenk et al. (2021).
@@
An increasing awareness of the importance of data quality has led to data filtering practices finding a more permanent place in the NLP pipeline, whereby Wikipedia’s encyclopaedic domain and culture of rigorous community moderation mean that it is typically assumed to be a ’high quality’ reference in this process (e.g., Wenzek et al., 2020). However, this assumption, while generally reliable for English and select higher-resourced languages, of Wikipedia as ‘high quality’, is now being challenged in the multilingual setting Kreutzer et al. (2022a). Wikipedias for lower-resourced languages can contain unnatural text written by non-native speakers, machine generated text, and other noise, which can be incomprehensible to native speakers Alabi et al. (2020); Lent et al. (2024). Poor data quality for low-resource languages can subsequently lead to poorly performing and untrustworthy models, ultimately resulting in unusable technologies for speakers Held et al. (2023); Nicholas and Bhatia (2023); Durmus et al. (2024).
@@
In this study, we critically examine the notion of Wikipedia as ‘high quality’ in a non-English setting. To this end, we compare common techniques for data filtering and how they perform on Wikipedia, particularly with attention to lower-resourced settings. We demonstrate the lack of apparent data quality in Wikipedia by showing that a lot of non-English Wikipedias contain foreign characters and duplicate articles, which we remove in a primary filtering step, and further apply corpus-driven heuristic quality filters. Downstream evaluation shows that these quality filters are not detrimental to performance, and thus can be used as a means for more resource-efficient training. Ultimately, we argue for a shift towards task and language specific approaches to data quality rather than seeking a general definition, especially when working with lower-resource languages. Our contributions are as follows:
@@
    We examine the effects of standard data cleaning techniques across non-English Wikipedias.
@@
    We demonstrate the importance of language and task specific approaches to data quality filtering.
With this, we add to the field’s larger debate about data quantity versus data quality. We aim for this study to be a starting guide to using Wikipedia as a resource for pretraining in a multilingual setting.
@@ 2.1 Data Filtering
Recent works in large language modelling have demonstrated the benefits of data filtering instead of pretraining over unpruned data dumps Wenzek et al. (2020); Raffel et al. (2023); Penedo et al. (2024); 01.AI et al. (2024). In this context, data filtering typically involves a combination of techniques, [^2] For a comprehensive survey of strategies for data filtering for LLMs, we refer readers to Albalak et al. (2024). ranging from simple heuristics and document deduplication to learned filters that estimate quality. Examples of heuristics from C4 include removing document containing code (e.g., Javascript), boiler plate legal policies (e.g., ‘Terms and Conditions’), or inappropriate terms, amongst other hand-crafted rules Raffel et al. (2023). Meanwhile, a popular technique for deduplication is MinHash, as it is easily scalable across CPU nodes Penedo et al. (2024); others combine MinHash with downstream clustering-based deduplication to filter more aggressively Tirumala et al. (2023). In addition to heuristics and deduplication, 01.AI et al. (2024) leverage learned filters (i.e., a quality classifier and cluster-based filters), which are designed to preserve documents resembling a ‘high quality’ reference corpus – and typically Wikipedia Wenzek et al. (2020); Together Computer (2023).
@@ 2.2 Wikipedia in Multilingual NLP
Wikipedia is the backbone of many massively multilingual benchmark datasets, such as WikiAnn Pan et al. (2017), XQuAD Artetxe et al. (2020), MLQA Lewis et al. (2020), TyDiQA-GoldP Clark et al. (2020), and BUCC Zweigenbaum et al. (2017), as well as a common source of pretraining data for popular multilingual LMs like mBERT Devlin et al. (2019) and XLM Lample and Conneau (2019). Even the newer and larger LMs (e.g., LLaMA Touvron et al. (2023a)), pretrained over enormous web corpora, can be critically dependent on Wikipedia. For example, CCNet’s language identification classifier was trained over the online encyclopedia Wenzek et al. (2020). Indeed, Wikipedia can be found in every part of the NLP pipeline from language identification to pretraining and fine-tuning, and conclusions drawn from massively multilingual studies are thus largely inextricable from Wikipedia Hu et al. (2020).
As a consequence of its ubiquity in NLP, Wikipedia is attracting increased scrutiny in terms of its data quality. A large scale quality audit of large web crawled datasets finds WikiMatrix Schwenk et al. (2021) has a number of deprecated, incorrect, or otherwise mislabelled language codes, among the 86 languages included Kreutzer et al. (2022a). Their additional manual audit of samples uncovers that quality gradually declines in correlation with resourcedness. Lignos et al. (2022) identify similar shortcomings for lower-resourced languages in WikiAnn and even Wikidata. [^3] https://wikidata.org Studies on selected low-resource languages further confirm this finding. For example, Alabi et al. (2020) observe that the Yorùbá and Twi Wikipedias are ostensibly written by non-native speakers and lack the necessary diacritics and variation to be relevant to speakers. Meanwhile, Lent et al. (2024) find that sentences from Wikipedias for Creole languages are often incomprehensible by speakers and require manual correction before being used to create benchmarks.
@@ 2.3 Data Quality versus Data Quantity
The notion of data quality is inherently subjective. Quantitatively, quality is typically measured by the relative boosting performance on a downstream task. Qualitatively, quality can be understood to mean samples which are accepted as valid by proficient speakers of a language, or samples that do not promote toxic or offensive sentiments Kreutzer et al. (2022b); van Noord et al. (2024). These definitions can sometimes be at odds with one another. For example, Longpre et al. (2024) show that a smaller, carefully filtered dataset can improve performance on downstream tasks in contrast to a larger, unfiltered dataset, but that the resulting model becomes more likely to generate toxic content. There are also diminishing gains from training on ever increasing sizes of data Warstadt et al. (2023). For pretraining LLMs, C4 (745GB) marginally outperforms Wikipedia (16GB) over benchmarks like GLUE, SGLEU, CNNDM, and SQuAD, as many of these benchmarks were sourced from Wikipedia, suggesting that quantity is less important than domain Raffel et al. (2023). For task-specific fine-tuning of large LMs, smaller, manually annotated datasets also lead to better performance than massive, open-source datasets Touvron et al. (2023b); Zhou et al. (2023); 01.AI et al. (2024); Gemini Team et al. (2024); Üstün et al. (2024). That said, data filtering is not a universally accepted practice; critics of filtering call attention to the risks of manipulating data, rather than using it ‘as it is’ Goldberg (2021). We refer readers to Rogers (2021) for a focused conversation on the benefits and drawbacks of data-curation within the context of training NLP models, both for general pre-training and for downstream tasks.
@@ 3 Is Wikipedia ‘High Quality’?
Albalak et al. (2024) outline a set of best practices for quality data selection and filtering. In this work, we examine the effects of their proposed model-agnostic filtering strategies [^4] We exclude classifier-based and perplexity-based filtering strategies measuring quality relative to a selection of data., as applied to Wikipedia. Conceptually, we can recognize such filters as belonging to two distinct categories based on their underlying objective: 1) procedures for removing undesirable material from Wikipedia, such as foreign scripts and duplicate articles; and 2) heuristics that aim to enhance the existing data distribution by removing otherwise lower quality text according to a given metric (e.g. article length in characters). For the first, we posit that if this step removes a substantial amount of text from a given Wikipedia dump, the removed content was not of high quality to begin with. We refer to this step as primary filtering. For the second, we aim to further refine Wikipedia articles via heuristic filters for even higher quality.
@@ 3.1 Primary Filtering
Applying an existing language identification model to Wikipedia is a challenge due to data contamination – most models supporting all of Wikipedia’s languages have generally also been trained on Wikipedia (see e.g. Kargaran et al., 2023). Thus, we employ script-based filtering as a proxy for language identification by filtering characters in the Wikipedia that are not in the officially documented script(s) [^5] For example, the Konkani Wikipedia has been documented to be written in 3 scripts – Devanagari, Latin and Kannada. We filter all characters that do not belong to any of these scripts. for the concerned language. [^6] https://en.wikipedia.org/wiki/List_of_Wikipedias Figure 1 demonstrates that 3.37% of characters in non-English Wikipedias are affected by this. Some Wikipedias even contain entire articles written in another language, generally English. [^7] We acknowledge that this methodology would not be effective for such English articles in non-English Wikipedias with Latin script. However, this would require a language ID model that has not been trained on Wikipedia (and therefore may have been trained on these articles as the ground truth.) For example, 6% of Assamese Wikipedia consists of foreign characters, where numerous articles are written almost entirely in English Wilson (2019). The implications of this are inevitably made clear when training a tokeniser on this Wikipedia, where much of the vocabulary budget can be used up by foreign characters.
@@
We employ hashing to identify and remove articles which have the exact same text, but a different unique ID or title. We find that some Wikipedias are strongly affected by this process. For example, the Yorùbá Wikipedia loses almost 60% of its articles. Typically, articles removed by exact-match deduplication are placeholders containing filler or templatic text and generally contain little usable information. Figure 1 shows that 1.01% of non-English Wikipedias articles are exact duplicates.
@@
While exact-match is able to filter identical documents, it cannot account for partial duplicates — i.e., sets of documents containing mostly identical text, but varying in terms of single sentences, paragraphs, or even hyperlinks. For example, the Yorùbá articles for the Baluch people and the Baluchi language are largely identical, with the latter containing an additional language data panel. To capture such cases, we perform minhash deduplication with locality sensitive hashing (LSH) and discount any article that returns a Jaccard similarity of 0.85. We find that a staggering 28.33% of all non-English Wikipedia articles (and 8.18% characters) are affected by this procedure — largely consisting of Wikipedias that have been known to be majority bot generated, such as Cebuano and Swedish. [^8] https://blog.datawrapper.de/wikipedia-articles-written-by-a-bot/
3.2 Heuristic Filtering
While the aforementioned primary filtering steps can reliably discount many undesirable documents, additional constraints may be considered in order to further refine the data selection process. To this end, we adopt what Albalak et al. (2024) categorise as heuristic filtering: a process in which document-level statistics are calculated in order to remove low quality data beyond a given threshold. Heuristics, in this sense, refer to simple metrics that aim to numerically characterise document content — for example, the number of characters contained in a given document. The red-pajama-data-v2 (Together Computer, 2023) dataset cleaning pipeline offers a collection of many such heuristics, culled from previous data curation efforts (Rae et al., 2021; Raffel et al., 2020; Longpre et al., 2024; Penedo et al., 2023). In theory, if Wikipedia is assumed to be a ‘high quality’ data source, it should be robust to these heuristics and not suffer from large-scale document removal.
To determine if this is the case, we explore natural language heuristics that are relevant to the encyclopedic nature of Wikipedia and easily applicable in a multilingual setting. [^9] For example, rps_doc_stop_word_fraction is not employed because it is difficult to principally obtain stopword lists for all the languages in Wikipedia. These are rps_doc_frac_unique_words (type-token ratio), rps_doc_unigram_entropy, and rps_doc_word_count — as collected by RedPajama. We also experiment with several adaptations to these metrics, such as utilising trigrams or characters as the basic unit of measurement (instead of whitespace-separated words). Naturally, some of these metrics demonstrate a very high correlation with each other; for example, absolute counts of characters, words, and trigrams typically exhibit a very high positive pairwise correlation across documents ($\rho>0.95$). As such, we decide to group all metrics into three classes: absolute, relative, and entropy. We list all metrics and their associated classes in Table 1.
@@ 3.2 Heuristic Filtering
Class Name Calculate…

### v2

@@ Abstract
Wikipedia’s perceived high quality and broad language coverage have established it as a fundamental resource in multilingual NLP. In the context of low-resource languages, however, these quality assumptions are increasingly being scrutinised. This paper critically examines the data quality of Wikipedia in a non-English setting by subjecting it to various quality filtering techniques, revealing widespread issues such as a high percentage of one-line articles and duplicate articles. We evaluate the downstream impact of quality filtering on Wikipedia and find that data quality pruning is an effective means for resource-efficient training without hurting performance, especially for low-resource languages. Moreover, we advocate for a shift in perspective from seeking a general definition of data quality towards a more language- and task-specific one. Ultimately, we aim for this study to serve as a guide to using Wikipedia for pretraining in a multilingual setting. We also introduce a package that implements our data cleaning, tokenization, and model training pipeline.
1 Introduction
@@
Wikipedia is an invaluable resource in NLP, with much of its usefulness steming from its open-source nature through which a community of authors, editors, and administrators can collaborate on the documentation of an ever-changing world. Such a framework ensures that information contained therein is not only current, but also well-written, accurate, and relevant to any given subject. For these reasons, Wikipedia has been employed for a myriad of NLP applications, such as language model (LM) training and benchmarking (Merity et al., 2016), question-answering (Hewlett et al., 2016; Rajpurkar et al., 2016; Joshi et al., 2017) and knowledge-base creation (Vrandečić and Krötzsch, 2014; Lehmann et al., 2024), among many others. Wikipedia is also a staple resource in multilingual NLP, where it is commonly employed for pretraining multilingual LMs (Devlin et al., 2019; Conneau and Lample, 2019) and creating benchmarks for downstream tasks, such as named-entity recognition Rahimi et al. (2019) and machine translation Schwenk et al. (2021).
@@
Recently, increasing attention to data quality has led to the adoption of various commonplace data filtering practices within NLP, whereby noisy web-scale data is distilled into a more salient signal for LM pretraining. Wikipedia has often been employed as a ‘high quality’ reference in this process (e.g., Wenzek et al., 2020), largely due to its encyclopaedic domain and culture of rigorous community moderation. However, the assumption of Wikipedia’s inherent ‘high quality’ — though generally reliable for English and select high-resource languages — is now increasingly being challenged in the multilingual setting Kreutzer et al. (2022a). Indeed, it has been found that Wikipedias for lower-resourced languages can contain unnatural, inorganic text written, often machine generated, and other noise, which can be incomprehensible to native speakers Alabi et al. (2020); Lent et al. (2024). As a result, poor data quality for such languages leads to degrading performance and untrustworthy models, ultimately providing unusable technologies for speakers Held et al. (2023); Nicholas and Bhatia (2023); Durmus et al. (2024).
@@
In this study, we critically examine the extent of Wikipedia’s ‘high quality’ in a non-English setting. Specifically, we compare common techniques for data filtering and assess how they perform on Wikipedia, with an emphasis on low-resource settings. We empirically demonstrate a lack of apparent data quality, showing that many non-English Wikipedias contain foreign characters and duplicate articles, which we remove in a primary filtering step, followed by the application of corpus-driven heuristic quality filters. Downstream evaluation reveals that our quality filters preserve, sometimes improve, downstream model performance, and thus can be used as a means for more resource-efficient training. Ultimately, we argue for a shift towards task and language-specific approaches to data quality, as opposed to one-size-fits-all approaches thereof — especially when working with lower-resource languages. Overall, our contributions are as follows:
@@
    We present empirical evidence of quality issues across Wikipedia and show that standard data cleaning pipelines can be used to mitigate them.
@@
    We make recommendations for using Wikipedia as a resource for dataset creation and pretraining models.
  4. 4.
    We compile a Wikipedia-specific data filtering pipeline and release code that can be used to replicate all our experiments. [^1] https://github.com/akulmizev/texieve.git
With this, we hope to provide a meaningful contribution to the field’s larger ongoing debate about data quantity versus quality. As such, we aim for this study to serve as a starting guide for using Wikipedia as a data resource in language modelling and beyond.
@@ 2.1 Data Filtering
Recent studies in large language modelling have demonstrated the benefits of filtering web-scale data Wenzek et al. (2020); Raffel et al. (2023); Penedo et al. (2024); 01.AI et al. (2024). In this context, data filtering typically involves a combination of techniques, ranging from basic document-level heuristics to corpus-level statistics that serve as quality proxies (Albalak et al., 2024). For example, the popular C4 dataset includes heuristics that remove documents containing code (e.g., ‘Javascript’), boiler plate legal policies (e.g., ‘Terms and Conditions’), or inappropriate terms, among other hand-crafted rules Raffel et al. (2023). Alongside heuristics, deduplication is another common filtering technique designed to remove duplicate documents from a given dataset, only keeping one copy thereof. A particularly popular algorithm is MinHash, due to its effectiveness in removing semantically identical documents, as well as its scalability across CPU nodes Penedo et al. (2024). MinHash has likewise been combined with downstream clustering-based deduplication to filter documents in a more aggressive manner Tirumala et al. (2023). It was also part of the data cleaning pipeline of the multilingual CulturaX 2024 dataset, in combination with language identification, URL-based filtering, and metric-based filtering with data-driven thresholds automatically identified. More sophisticated techniques have recently also been employed: for example, 01.AI et al. (2024) leverage learned filters (i.e., quality classifiers and cluster-based filters), which are designed to preserve documents resembling a ‘high quality’ reference corpus – typically Wikipedia Wenzek et al. (2020); Together Computer (2023).
@@ 2.2 Wikipedia in Multilingual NLP
Wikipedia’s widespread language coverage and reputation as a high-quality data trove has made it a vital resource in multilingual NLP applications. It has served as the backbone of many multilingual benchmark datasets, such as WikiAnn Pan et al. (2017), XQuAD Artetxe et al. (2020), MLQA Lewis et al. (2020), TyDiQA-GoldP Clark et al. (2020), and BUCC Zweigenbaum et al. (2017), as well as a common source of pretraining data for popular multilingual LMs, such as mBERT Devlin et al. (2019) and XLM Lample and Conneau (2019). Beyond this, it was notably employed as the training data for CCNet’s language identification classifier, which is primarily used for partitioning the popular CommonCrawl dataset along language boundaries. CommonCrawl serves as the basis for many contemporary open-source pretraining datasets, therefore the quality and performance of such a classifier is paramount in retaining relevant data.
As a consequence of its ubiquity in NLP, Wikipedia has recently attracted increased scrutiny for its data quality. For example, Kreutzer et al. (2022a) find that the Wikipedia-oriented WikiMatrix Schwenk et al. (2021) contains a number of deprecated, incorrect, or otherwise mislabelled language codes, among the 86 languages included Kreutzer et al. (2022a). Lignos et al. (2022) identify similar shortcomings for lower-resourced languages in WikiAnn and even Wikidata. [^2] https://wikidata.org Studies on selected low-resource languages likewise confirm such findings. For example, Alabi et al. (2020) observe that the Yorùbá and Twi Wikipedias are ostensibly written by non-native speakers and lack the necessary diacritics and variation to be relevant to speakers. Meanwhile, Lent et al. (2024) find that sentences from Wikipedias for Creole languages are often incomprehensible by speakers and require manual correction before being used to create benchmarks.
@@ 2.3 Data Quality versus Data Quantity
The notion of data quality is inherently subjective. Quantitatively, it is commonly measured by performance gains on a downstream NLP task. Qualitatively, it can be understood to signify utterances that are accepted as valid by proficient speakers of a language, or, alternatively, text samples that do not promote toxic or offensive sentiments Kreutzer et al. (2022b); van Noord et al. (2024). Such definitions can sometimes be at odds with one another, however. For example, Longpre et al. (2024) show that a smaller, carefully filtered dataset can improve performance on downstream tasks in contrast to a larger, unfiltered dataset, but that the resulting model becomes more likely to generate toxic content. Likewise, diminishing gains have been demonstrated from training on ever-increasing amounts of data Warstadt et al. (2023). For example, language models pretrained on C4 (745GB) marginally outperform those trained on Wikipedia (16GB) over benchmarks like GLUE (Wang et al., 2018), SuperGLUE (Wang et al., 2019), and SQuAD (Rajpurkar et al., 2016). However, many of these benchmarks were sourced from Wikipedia itself, suggesting that quantity is less important than domain Raffel et al. (2023).
@@ 3 Is Wikipedia ‘High Quality’?
Albalak et al. (2024) outline a set of best practices for quality data selection and filtering. In this work, we examine the effects of their proposed model-agnostic filtering strategies, [^3] We exclude classifier-based and perplexity-based filtering strategies measuring quality relative to a selection of data. as applied to Wikipedia. Conceptually, we can recognize such filters as belonging to two distinct categories based on their underlying objective: 1) procedures for removing undesirable material from Wikipedia, such as foreign scripts and duplicate articles; and 2) heuristics that aim to enhance the existing data distribution by removing otherwise lower quality text according to a given metric (e.g. article length in characters). For the first category, we posit that if this step removes a substantial amount of text from a given Wikipedia dump, the removed content was not of high quality to begin with. We refer to this step as primary filtering. For the second, we aim to further refine Wikipedia articles via heuristic filters for even higher quality. Such a categorisation is also employed in the creation of MADLAD-400 Kudugunta et al. (2023), where text is deduplicated in a preliminary step, followed by language identification and then heuristic-based filters are applied.
@@ 3.1 Primary Filtering
Applying an existing language identification model to Wikipedia is a challenge due to data contamination – most models supporting all of Wikipedia’s languages have generally also been trained on Wikipedia (see e.g. Kargaran et al., 2023). Thus, we employ script-based filtering as a proxy for language identification by filtering characters in the Wikipedia that are not in the officially documented script(s) [^4] For example, the Konkani Wikipedia features 3 scripts – Devanagari, Latin and Kannada. for the concerned language. [^5] https://en.wikipedia.org/wiki/List_of_Wikipedias Figure 1 demonstrates that 3.37% of characters in non-English Wikipedias are affected by this. Additionally, various Wikipedias contain full articles written in a different language, generally English. [^6] We acknowledge that this methodology would not be effective for such English articles in non-English Wikipedias written in the Latin script. However, this would require a language ID model that has not been trained on Wikipedia For example, 6% of Assamese Wikipedia consists of foreign characters, where numerous articles are written almost entirely in English Wilson (2019). The implications of this are inevitably made clear when training a tokeniser on this Wikipedia, where much of the vocabulary budget can be used up by foreign characters.
@@
We employ a hashing function to identify and remove articles which have the exact same text, but a different unique ID or title. We find that some Wikipedias are strongly affected by this process — for example, the Yorùbá Wikipedia loses almost 60% of its articles — many of which contain a single token Ìtọḱasí (Reference). Typically, articles removed by exact-match deduplication are placeholders containing filler or templatic text and generally contain little usable information. Figure 1 shows that 1.01% of non-English Wikipedias articles are exact duplicates.
@@
While exact-match is able to filter identical documents, it cannot account for partial duplicates — i.e., sets of documents containing mostly identical text, but varying in terms of single sentences, paragraphs, or even hyperlinks. For example, the Yorùbá articles for the Baluch people and the Baluchi language are largely identical, with the latter containing an additional language data panel. To capture such cases, we perform MinHash deduplication with locality sensitive hashing (LSH) and discount any article that returns a Jaccard similarity of 0.85. We find that a staggering 28.33% of all non-English Wikipedia articles (and 8.18% characters) are affected by this procedure, largely consisting of Wikipedias known to be majority bot generated (more in §4.3).
@@
Class Name Calculation
@@
3.2 Heuristic Filtering
Albalak et al. (2024) categorise a process in which document-level statistics are calculated in order to remove low quality data beyond a given threshold as heuristic filtering. Heuristics, in this sense, refer to simple metrics that aim to numerically characterise document content — for example, the number of characters contained in a given document. The red-pajama-data-v2 (Together Computer, 2023) dataset cleaning pipeline offers a collection of many such heuristics, culled from previous data curation efforts (Rae et al., 2021; Raffel et al., 2020; Longpre et al., 2024; Penedo et al., 2024). In theory, if Wikipedia is assumed to be a ‘high quality’ data source, it should be robust to these heuristics and not suffer from large-scale document removal.
To determine if this is the case, we explore natural language heuristics that are relevant to the encyclopedic nature of Wikipedia and easily applicable in a multilingual setting. [^7] For example, rps_doc_stop_word_fraction is not employed because it is difficult to principally obtain stopword lists for all the languages in Wikipedia. These are rps_doc_frac_unique_words (type-token ratio), rps_doc_unigram_entropy, and rps_doc_word_count – as collected by RedPajama. We also experiment with several adaptations to these metrics, such as utilising trigrams or characters as the basic unit of measurement (instead of whitespace-separated words). Naturally, some of these metrics demonstrate a very high correlation with each other; for example, absolute counts of characters, words, and trigrams typically exhibit a very high positive pairwise correlation across documents ($\rho>0.95$). As such, we decide to group all metrics into three classes: absolute, relative, and entropy. We list all metrics and their associated classes in Table 1.